For z€c, minimum value of |z-4+3i|+|z+3-4i|is A) 10B) 0C) 2(7)^1/2D) 2

Grade:11

1 Answers

ASHWIN SRIKANTH
7 years ago
​generally  minimum value of \left | z1-z2 \right | is \left | z2 \right |​−\left | z1 \right |   and  \left | z1+z2\right | = \left | z2 \right |​+\left | z1 \right | applying the same here we get the 
5-z+z+5 = 10 
 
as \left | 3+4i\right |= 5 therefore the required answer is A.thank you